8/5 x (-6/5) x 8/5
Answer by Alan3354(69443)

8/5 x (-6/5) x 8/5
---------
Use * (Shift 8) for multiply
8/5*(-6/5)*8/5 = (8*(-6)*8)/(5*5*5) = -384/125